OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5).

OpenStudy (anonymous):

So you know the gradient that you are looking for is (was it -5?) the equation you then form is y = mx+b in this case m = -5 so sub x = 2 and y = 5 to the equation to find b and re-write the answer.

OpenStudy (anonymous):

first find the slope of the given line slope of the given line =- (coefficient of x)/(coefficient of y) = m (say) then the slope of the line perpendicular to the given is given by -1/m now eqyuation of line with slope m1 and passing throughx1,y1 is given by y-y1= m1 (x-x1)

OpenStudy (anonymous):

here slope of the given line =1/5 hnce slope of the line perpendicular to it is given by -5 and since the line passes through 2,5 and hence the reqd line is (y-5) = -5 (x-2)

OpenStudy (anonymous):

y = 2x ?

OpenStudy (anonymous):

You set the problem in slope-intercept form: y=mx+b. Then you take the slope, m, and make it into the negative reciprocal. With that, use the formula, y-y1=m2(x-x1), and plug in the two points for y1 and x1. Then plug in the new slope, and distribute it, then solve for y. You should get y=-5x+5, and I'm not sure whether or not you need to find b.

OpenStudy (anonymous):

these are the questions Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). and these are the answers A) y = -2x - 1 B) y = -3x + 14 C) y = 2x D) y = -5x + 15 E) y = -4x + 20